Zonenplatte

Golubkov Andrej

Meierhofer Georg Betreuer: Dr. Bodil Holst

10. Mai 2004

Inhaltsverzeichnis

1 Tabelle 3 2 Graphik einfügen 3

1 Tabelle

132  75 8,2 32525
2432  12400 8,4 68
54323 443225 8,6 172

2 Graphik einfügen

Dieses Kapitel enthält die Graphik Abb.??.